OpenStudy (anonymous):

Can someone please explain to me how to find the Fourier Transform of sinc(2t)?

OpenStudy (experimentx):

what is sinc(t) sine integral?

OpenStudy (anonymous):

I don't really know :/

OpenStudy (anonymous):

I know sinc(t) = \[\sin (\pi t)/\pi t\]

OpenStudy (experimentx):

use Dirichlet integral

OpenStudy (experimentx):

http://en.wikipedia.org/wiki/Dirichlet_integral

OpenStudy (anonymous):

I still don't understand... the answer is apparently \[(\pi/2) rect(\omega/4)\]

OpenStudy (anonymous):

I don't know how they obtained it, I tried using the formula.

OpenStudy (experimentx):

http://www.wolframalpha.com/input/?i=Fourier+Transform+Sinc%5BPi+t%5D \[ \begin{align*} f(\omega) &= \frac{1}{\sqrt{2 \pi }} \int_{-\infty}^\infty \frac{\sin (\pi t)}{(\pi t)} e^{-i \omega t}dt \\ &= \frac{1}{\sqrt{2 \pi }} \left( \int_{-\infty}^0 \frac{\sin (\pi t)}{(\pi t)} e^{-i \omega t}dt + \int_0^{\infty} \frac{\sin (\pi t)}{(\pi t)} e^{-i \omega t}dt\right )\\ &= \frac{1}{\sqrt{2 \pi }} \left( \int_0^{\infty} \frac{\sin (\pi t)}{(\pi t)} e^{i \omega t}dt + \int_0^{\infty} \frac{\sin (\pi t)}{(\pi t)} e^{-i \omega t}dt\right )\\ &= \frac{1}{\sqrt{2 \pi }} \int_0^{\infty} \frac{\sin (\pi t)}{(\pi t)} (e^{i \omega t} +e^{i \omega t}) dt\\ &= \frac{1}{\sqrt{2 \pi }} \int_0^{\infty} \frac{\sin (\pi t)}{(\pi t)} (2 \cos(\omega t)) dt \\ &= \frac{2}{\sqrt{2 \pi }}\int_0^{\infty} \frac{\sin (\pi t)}{(\pi t)} (\cos(\omega t)) dt \end{align*} \] You can plug this integral into WA and solve for it. There are many ways you can evaluate this integral

OpenStudy (experimentx):

\[ \begin{align*} \int_0^{\infty} \frac{\sin (\pi t)}{(\pi t)} \cos(\omega t) dt &= \int_0^\infty \frac{\sin ((\pi + \omega)t) +\sin ((\pi - \omega)t) }{\pi t} dt\\ &= \int_0^\infty \frac{\sin ((\pi + \omega)t)}{ \pi t}dt + \int_0^\infty \frac{\sin ((\pi - \omega)t)}{ \pi t}dt\\ \end{align*} \] \[ \int_0^\infty \frac{\sin (kx)}{x}dx = \text{sign}(k) \frac \pi 2 \]